OpenStudy (anonymous):

If you write the function P=10e^−2t in the form P=P0a^t, then what is P0 and a?

OpenStudy (anonymous):

i know P0 is equal to 10 but not sure what a is equal to

OpenStudy (anonymous):

this also shows decay

OpenStudy (phi):

the trick is to use this idea with exponents \[ \left(a^b\right)^c = a^{bc} \]

OpenStudy (phi):

or I should say, the other way round: \[ a^{bc}=\left(a^b\right)^c \] notice we can match e with the "a" ,-2 with b, and "t" with c in \[ e^{−2t} \] can you rewrite this to show the "t" "pulled out" like in \[ a^{bc}=\left(a^b\right)^c \]?

OpenStudy (phi):

if it doesn't show, refresh your browser.

OpenStudy (anonymous):

(a^-2)^t @phi ?

OpenStudy (phi):

almost. it is still "e" (e^-2)^t the the stuff in the parens e^-2 is a number. Use a calculator or type e^-2 = into google

OpenStudy (phi):

in other words \[ 10 e^{-2t} = 10 \ \left( e^{-2}\right)^t \\ = 10 (0.13533)^t \]

OpenStudy (phi):

roughly... e^-2 goes on forever, so we have to stop at some point. it's now in this form P=P0 a^t and as you already know P0 = 10 and a = e^-2 = (about) 0.13533...

OpenStudy (loser66):